# SMARTER: A Data‑efficient Framework to Improve Toxicity Detection with Explanation via Self-augmenting Large Language Models

### v2

@@ Abstract
WARNING: This paper contains examples of offensive materials. To address the proliferation of toxic content on social media, we introduce SMARTER, a data-efficient 2-stage framework for explainable content moderation using Large Language Models (LLMs). In Stage 1, we leverage LLMs’ own outputs to generate synthetic explanations for both correct and incorrect labels, enabling alignment via preference optimization with minimal human supervision. In Stage 2, we refine explanation quality through cross-model training, allowing weaker models to align stylistically and semantically with stronger ones. Experiments on 3 benchmark tasks (HateXplain, Latent Hate, and Implicit Hate) demonstrate that SMARTER enables LLMs to achieve up to 13.5% macro-F1 improvement over standard few-shot baselines with a fraction of the full training data. Our framework offers a scalable strategy for low-resource settings by harnessing LLMs’ self-improving capabilities for both classification and explanation.
@@ 1 Introduction
Toxic content encompasses a wide spectrum of terminologies whose definitions vary by platform: hate speech, cyberbully, sexist, racist etc. Gelber (2021); Anjum and Katarya (2024). Typically, human moderators must manually adjudicate the target content, an inefficient process taxing to the moderators themselves Baker et al. (2020); Spence et al. (2023). Recently, social media platforms can significantly reduce this overhead by integrating machine learning models to automate the detection pipeline d’Sa et al. (2020); Sumanth et al. (2022).
@@
    We introduce SMARTER [^1] Our repository can be found at REDACTED (Self-augMentAtion Regimen Towards Efficient Content ModeRation), a 2-stage framework where (1) each LLM self-augments on few-shot data using preference optimization with synthetic explanations, (2) cross-model refinement to further improvement by training one LLM with responses generated by another on held-out sets.
@@
    Experiments conducted on 3 benchmark datasets (HateXplain, Latent Hate, Implicit Hate) with Llama-3.1-8B-Instruct and COT-T5-XL demonstrate up to 13.5% increase in macro-F1 scores in few-shot settings with high-quality explanations, outperforming a range of commercial models.
SMARTER thus provides a scalable and practical approach for low-resource NLP practitioners to achieve strong performance and explainability.
2 Related Works
@@
Existing research Caselli et al. (2021); Sarkar et al. (2021); Toraman et al. (2022) has proposed strategies to enhance cross-domain robustness in toxicity detection, such as combining multiple datasets to improve generalization Antypas and Camacho-Collados (2023) and incorporating label definitions to support few-shot adaptation Nghiem et al. (2024). Building on these advances, recent studies leverage LLMs for toxicity detection, demonstrating strong zero- and few-shot performance Masud et al. (2024); Jahan et al. (2024); Roy et al. (2023); Kumarage et al. (2024). LLMs can also generate textual explanations, enhancing interpretability for moderators and users Di Bonaventura et al. (2024); Yang et al. (2023). We extend this line of work by training open-source LLMs to jointly perform classification and explanation to promote transparency in moderation workflows.
@@
In addition to extensive pretraining, modern LLMs also undergo post-training that both unlocks additional capabilities Longpre et al. (2023) and aligns the outputs towards human preferences Ouyang et al. (2022). Recent Reinforcement Learning (RL)-based techniques like GRPO Shao et al. (2024); Zweiger et al. (2025) enable promising LLM self-alignment at expensive computational cost. Another body of work showcases LLMs to self-improve via their own synthetic data Tang et al. (2023); Li et al. (2023), such as Self-Refine Madaan et al. (2023) and Self-Instruct Wang et al. (2022). Building upon these ideas, our framework presents a simple yet principled approach that allows LLMs to iteratively improve with minimal training cost while still amenable to human supervision.
@@ 4.1 Data
We select 2 benchmark datasets that cover distinct yet still relevant concepts to toxic content moderation and derive from them 3 classification tasks.
@@
Definitions of the associated labels are also included, allowing us to integrate these contexts into the explanation in downstream applications in a fashion similar to Yang et al. (2023) and Nghiem and Daumé (2024). We refer to the three classification tasks by their label spaces as HateXplain, Latent Hate, and Implicit Hate. Table 4 summarizes their components and data splits, with preprocessing details in Appendix B.
@@ 4.3 Classification via Prompting
We apply this Chain-of-Thought (COT) paradigm Wei et al. (2022) in our classification tasks to obtain the explanation from the LLMs necessary for content moderation. Inspired by Nghiem and Daumé (2024) and Yang et al. (2023), we use the prompt in Figure 6, which incorporates the target post, the set of labels corresponding to the task and their associated definitions as context for the LLM to provide its explanation and predicted category.
@@ 5.1 K-shot In-domain Finetuning
We experiment with different configurations of limited in-domain data. For each task, we select uniformly at random K-shot $\in\{16,32,64,128,256\}$ samples from the training set . We utilize the seed explanations by Nghiem and Daumé (2024) to further SFT the existing HateCOT-pretrained T5 and Llama models in a fashion similar to the auxiliary pretraining phase above.
@@ 5.1.1 Experimental Pipeline
The experiments follow a two-stage process. First, we finetune T5 and Llama under extreme data scarcity by setting $K\in\{16,32,64,128\}$, followed by self-augmentation via DPO to compare classification performance. Then, we experiment with significantly more training data with $K=256$ and evaluate KTO alignment against DPO  [^2] Implementation for KTO trainer for T5 is not available at the time of writing, so we are only able to perform this variant for Llama. For comparison, we also finetune both LLMs with the entire available training data to obtain classification label without explanation, denoted Full models. In addition, we also finetune ModernBERT (large) Warner et al. (2024), a decoder-only model on the full training data and obtain the classification outputs using the same test sets.
@@ 5.2.1 Results for $K<=128$
We report macro-F1–the harmonic mean between Precision and Recall–as the evaluation metric Christen et al. (2023). Figure 1 displays the bar plots of the LLM’s Baseline and DPO-enhanced performance on the test sets.
Self-augmentation with DPO generally boosts classification performance over its Baseline counterparts, with the sole reversal for T5 on Latent Hate at $K=16$ (2.1% reduction) along with static scores for T5 on HateXPlain at $K=32$, and for Llama in Latent Hate at $K=64$. The infrequency of these observations and the negligible decrease highlight our method’s consistent enhancement. In fact, increase in F1 scores go up to 14.5% for Llama (HateXplain, $K=128$) and 25% for T5 (Implicit Hate, $K=64$).
@@
Figure 1: Bar plots for $K$-shot classification experiments on 3 datasets using Llama and T5 models. Macro F1 scores and percentage change over Baseline are displayed on top. Results for Baselines and DPO-augmented variants for $K\in\{16,32,64,128\}$ are displayed on the left subfigures. Results for $K=256$ of Baseline, KTO, DPO-augmented and its other variants on various sub-sampling strategies (section 5.1.1) are shown on the right. Horizontal lines show the F1 scores for Full models that use all training data.
@@ 5.2.2 Results for $K=256$
In Figure 1, we observe the macro F1 scores for models $K=256$ on the subfigures on the right. DPO-augmentation continues to enhance Baseline models which already benefits from more training data relative to lower $K$ values, with improvement in scores reaching up to 13.2% for LLama and 14% for T5. In contrast, KTO-augmentation appears ineffectual on HateXplain and even hinders performance on Latent Hate and Implicit Hate, showing notable reduction from Baseline scores, respectively. With this, we hypothesize that DPO’s contrastive nature reinforces signals that allow the LLM to learn from its own mistakes, thereby sharpening its reasoning ability for this type of discriminatory task. On the other hand, KTO’s reliance on implicit signals from listwise data is not sufficient to improve discernment of classes when the semantic distinction between labels are nuanced for Latent Hate and Implicit Hate.
@@
Dataset Model F1 F1% Data%
HateXplain Llama_DPO 0.64 89% 6%
Llama_Full 0.72 100% 100%
T5_DPO 0.62 86% 6%
T5_Full 0.72 100% 100%
ModernBERT 0.70 97% 100%
GPT-4o-mini 0.50 69%
GPT-4.1 0.55 76%
GPT-5-chat 0.56 78%
Qwen-32B 0.54 75%
Latent Hate Llama_DPO 0.69 100% 7%
Llama_Full 0.62 90% 100%
T5_DPO 0.65 94% 7%
T5_Full 0.61 88% 100%
ModernBERT 0.61 88% 100%
GPT-4o-mini 0.54 78%
GPT-4.1 0.60 87%
GPT-5-chat 0.51 74%
Qwen-32B 0.47 68%
Implicit Hate Llama_DPO 0.60 90% 57%
Llama_Full 0.67 100% 100%
T5_DPO 0.59 88% 57%
T5_Full 0.64 96% 100%
ModernBERT 0.64 96% 100%
GPT-4o-mini 0.42 63%
GPT-4.1 0.57 85%
GPT-5-chat 0.58 87%
Qwen-32B 0.49 73%
Table 1: Comparison of performance and data usage across models (including additional GPT baselines). We report Macro F1 (F1), relative score (F1%), and training data fraction (Data%). Models with the DPO suffix are DPO-augmented with K=256K{=}256 shots, while Full uses the entire training set. Within each dataset, F1 is bolded for the best model, and F1% is normalized to that best model (100%). Data% is measured relative to the corresponding Full baseline (100%). ModernBERT is shown for reference.
Comparison against Alternative Baselines
We further evaluate zero-shot classification on three commercial OpenAI LLMs and Qwen-32B, alongside off-the-shelf T5 and Llama variants, using the prompt in Figure 5. As shown in Table 1 and Table 5, our DPO-augmented models at $K=256$ consistently outperform both their baselines and larger LLMs, demonstrating SMARTER’s accessibility advantage—achieving controllable, high-quality performance with lightweight, low-cost models that do not rely on proprietary platforms.
Following the same setup as ModernBERT, we also SFT-finetune base Llama and T5 for classification without explanations. At $K=256$, our DPO-augmented variants retain at least 86% of the Full model’s F1 on HateXplain and surpass it on Latent Hate, underscoring the efficiency and scalability of our low-resource alignment framework.
@@ 6 Stage 2: Cross-model Refinement
In this section, we focus on investigating methods to identify and efficiently improve the LLMs’ quality of explanations with respect to certain desired directions. We first enlist human annotators to evaluate the explanations by T5 and Llama, then experiment with further training one model with the other’s outputs to simultaneously improve performance and quality of explanations.
6.1 Human Evaluation
To simulate real-life deployment, we solicit 14 crowdworkers from diverse demographic backgrounds on the platform Amazon Mechanical Turks (MTurks) to annotate their preference on the explanations offered by the LLMs in this study . Annotators are encouraged to evaluate the pairs of explanation for each post based on the criteria of Clarity, Reasoning and Alignment as illustrated with the template in Figure 3. To avoid position bias, we present the pairs of explanations in randomized orders. All annotators are compensated fairly.
@@ 6.1 Human Evaluation
Table 2: Results of preference for explanations by MTurk annotators based on majority voting for 342 HateXplain samples.
@@ 6.2.1 Cross-model Training Methodology
Since Llama produces outputs preferred by human annotators, a natural choice is to select this model for ultimate deployment. We are nevertheless interested in examining the potential to steer one model’s explanation style to align with another model’s while maintaining existing classification performance.
Using the prompting technique described in section 6.1, we use the SFT+DPO $K=128$ variants of both Llama and T5 LLMs to generate explanations conditioned on the gold labels of the held-out 128-shot portion of the training data, which neither of these models has seen. We then perform cross-model SFT and DPO finetune using this dataset. More specifically, we further finetune the existing SFT+DPO $K=128$ T5 variant using explanations produced by its Llama counterpart, and vice versa. This step consists of both separate SFT and then DPO self-augmentation as in section 5.1.
@@ 6.2.3 Cross-model Style Analysis
From Figure 11, most explanations from the individually trained models exhibit their corresponding style, while cross-model SFT is sufficient to substantially align the outputs with the alternative’s style. With the exception of T5 on HateXplain, the application of self-augmenting DPO using cross-model data generally amplifies the stylistic presence of the base model to a small extent, suggesting that this mechanism induces a mild converging effect on the reasoning style.
@@ 7 Analysis of Explanations
We perform Natural Language Inference (NLI) evaluation to ascertain whether the LLMs’ explanations remain internally consistent with their corresponding labels and definitions throughout the augmentation pipeline, and then anatomize the stylistic changes between cross-model refinement.
@@ 7.1 Analysis of Consistency
We prompt GPT-4o-mini with the template shown in Figure 8 to judge whether the offered explanations is logically consistent with the predicted label for each sample across 3 test sets via the 4 categories: Entail, Contradict, Neutral and Undefined (noise cases where either the predicted label or the explanation is missing). An author’s independent annotation of 150 random samples from this pool under the same instruction achieves nearly unanimous agreement with GPT-4o-mini (Cohen’s $\kappa=0.92$).
@@ 8 Discussion and Conclusion
Informed by empirical insights , we offer the following recommendations for practitioners.
Cross-model refinement can further improve performance
T5’s superior F1-scores compared to individually trained models (Figure 2) suggest added value in leveraging competing models. Practitioners can train multiple LLMs using the SFT+DPO framework across data thresholds, then select a high-quality checkpoint (e.g., $K=128$ from Section 5.2.1) for further finetuning with outputs from other models. Although further gains are not guaranteed, cross-model refinement represents a practical and resource-efficient direction for improving alignment and robustness.
Architectural diversity may offer benefits.
The differing post-training behaviors of T5 and Llama—despite being trained on the same seed explanations—may stem from their distinct architectures (encoder-decoder vs. encoder-only). We hypothesize based on prior work Hao et al. , that the stronger model (Llama) enhances reasoning diversity that T5 can leverage. While further study is needed for a definitive explanation, our findings suggest practitioners may benefit from using architecturally diverse models within this framework.
Human oversight is still essential.
Beyond the initial annotation, human supervision is necessary to monitor the stylistic quality of explanations, especially in iterative-development settings where unexpected patterns may occur after multiple rounds of cross-model training.

### v3

@@ Abstract
WARNING: This paper contains examples of offensive materials. To address toxic content on social media, we introduce SMARTER, a data-efficient 2-stage framework for explainable content moderation using Large Language Models (LLMs). In Stage 1, we leverage LLMs’ own outputs to generate synthetic explanations for correct and incorrect labels, enabling preference optimization with minimal supervision. In Stage 2, we refine explanation quality through cross-model training, allowing weaker models to align with stronger ones. Experiments on 3 classification tasks (HateXplain, Latent Hate, Implicit Hate) show SMARTER achieves up to 13% macro-F1 improvement over few-shot baselines using only 6-57% of training data. Our framework offers a scalable strategy for low-data settings by harnessing LLMs’ self-improvement for explainable moderation.
@@ 1 Introduction
Toxic content encompasses a wide spectrum of terminologies whose definitions vary by platform: hate speech, cyberbully, sexist, racist etc. Gelber (2021); Anjum and Katarya (2024); Wang et al. (2025). Typically, human moderators must manually adjudicate the target content, an inefficient process taxing to the moderators themselves Baker et al. (2020); Spence et al. (2023). Recently, social media platforms can significantly reduce this overhead by integrating machine learning models to automate the detection pipeline d’Sa et al. (2020); Sumanth et al. (2022).
@@
    We introduce SMARTER [^1] Our repository can be found at https://github.com/hnghiem-nlp/hate_dpo_public (Self-augMentAtion Regimen Towards Efficient Content ModeRation), a 2-stage framework where (1) each LLM self-augments on few-shot data using preference optimization with synthetic explanations, (2) cross-model refinement to further improvement by training one LLM with responses generated by another on held-out sets.
@@
    Experiments conducted on 3 classification tasks (HateXplain, Latent Hate, Implicit Hate) with Llama-3.1-8B-Instruct and COT-T5-XL demonstrate up to 13.5% increase in macro-F1 scores in few-shot settings with high-quality explanations, outperforming a range of commercial models.
While commercial APIs may incur high costs and limited transparency, SMARTER delivers competitive performance and explainability with 6-57% of training data, ensuring deployment control.
2 Related Work
@@
Existing research Caselli et al. (2021); Sarkar et al. (2021); Toraman et al. (2022) has proposed strategies to enhance cross-domain robustness in toxicity detection, such as combining multiple datasets to improve generalization Antypas and Camacho-Collados (2023) and incorporating label definitions to support few-shot adaptation Nghiem et al. (2024). Building on these advances, recent studies leverage LLMs for toxicity detection, demonstrating strong zero- and few-shot performance Masud et al. (2024); Jahan et al. (2024); Roy et al. (2023); Kumarage et al. (2024). LLMs can also generate textual explanations, enhancing interpretability for moderators and users Di Bonaventura et al. (2024); Yang et al. (2023); Calabrese et al. (2024); Almohaimeed et al. (2025). We extend this line of work by training open-source LLMs to jointly perform classification and explanation to promote transparency.
@@
In addition to extensive pretraining, modern LLMs also undergo post-training that both unlocks additional capabilities Longpre et al. (2023) and aligns the outputs towards human preferences Ouyang et al. (2022). Recent Reinforcement Learning (RL)-based techniques Shao et al. (2024); Zweiger et al. (2025); Nghiem et al. (2025) enable promising LLM self-alignment at expensive computational cost. Another body of work showcases LLMs to self-improve via their own synthetic data Tang et al. (2023); Li et al. (2023), such as Self-Refine Madaan et al. (2023) and Self-Instruct Wang et al. (2022). Building on these ideas, our framework presents a simple yet principled approach that allows LLMs to iteratively improve with minimal training cost while amenable to human supervision.
@@ 3 Overview of Framework
Input: It means people aren’t buying the lie! diversity k**s now prove me wrong! Explanation: The post uses coded or indirect language to disparage a person or group based on their race. Overall, the use of the phrase "diversity k**s" and the challenge to "prove me wrong" indicate a negative view of diversity, suggesting a belief that diversity is harmful. This coded language indirectly disparages people of color, making it an example of implicit hate. Label: Implicit Hate
Figure 2: Example of explanation and classification label produced by Llama at $K=128$ shot on the dataset Latent Hate. For definition of categories, see Table 10.
@@ 4.1 Data
We select 2 datasets that cover distinct yet still relevant concepts to toxic content moderation and derive from them 3 classification tasks.
@@
Label definitions are included, enabling context integration into explanations following Yang et al. (2023) and Nghiem and Daumé (2024). We denote these three tasks as HateXplain, Latent Hate, and Implicit Hate. Table 4 summarizes data splits; preprocessing details are in Appendix B. Figure 2 provides an example of the explanation produced by our model on Latent Hate.
@@ 4.3 Classification via Prompting
We apply the Chain-of-Thought (COT) paradigm Wei et al. (2022) in our classification tasks to obtain the explanation from the LLMs necessary for content moderation. Inspired by Nghiem and Daumé (2024) and Yang et al. (2023), we use the prompt in Figure 7, which incorporates the target post, the set of labels corresponding to the task and their associated definitions as context for the LLM to provide its explanation and predicted category.
@@ 5 Stage 1: Individual Model Training
We demonstrate SMARTER’s data efficiency by training models at extreme low-shot settings and evaluating against off-the-shelf, commercial, and full-training baselines.
@@ 5.1 K-shot In-domain Finetuning
For each task, we select uniformly at random K-shot without replacement $\in\{16,32,64,128,256\}$ samples from the training set . We utilize the seed explanations by Nghiem and Daumé (2024) to further SFT the existing HateCOT-pretrained T5 and Llama models similar to the auxiliary pretraining phase above.
@@ 5.1.1 Experimental Pipeline
The experiments follow a two-stage process. We first finetune T5 and Llama at $K\in\{16,32,64,128\}$ followed by DPO self-augmentation. At $K=256$, we compare KTO and DPO alignment. [^2] KTO trainer for T5 is unavailable, so we evaluate KTO only for Llama. For comparison, we finetune both LLMs on full training data for classification without explanations (Full models), and similarly train ModernBERT (large) Warner et al. (2024) as an additional baseline.
@@ 5.2.1 Results for $K<=128$
We report macro-F1–unweighted mean of per-class F1 scores–as the evaluation metric Christen et al. (2023). Figure 3 displays the bar plots of the LLM’s Baseline and DPO-enhanced performance on the test sets.
Self-augmentation with DPO generally boosts classification performance over Baseline counterparts, with the sole reversal for T5 on Latent Hate at $K=16$ (2.1% reduction), along with static scores for T5 on HateXplain at $K=32$ and Llama on Latent Hate at $K=64$. These infrequent and negligible exceptions highlight our method’s consistent enhancement. For instance, F1 gains reach 14.5% for Llama (HateXplain, $K=128$) and 25% for T5 (Implicit Hate, $K=64$).
@@
Figure 3: Bar plots for $K$-shot classification experiments on 3 datasets using Llama and T5 models. Macro-F1 scores and percentage change over Baseline are displayed on top. Results for Baselines and DPO-augmented variants for $K\in\{16,32,64,128\}$ are displayed on the left subfigures. Results for $K=256$ of Baseline, KTO, DPO-augmented and its other variants on various sub-sampling strategies (section 5.1.1) are shown on the right. Horizontal lines show the F1 scores for Full models that use all training data.
@@ 5.2.2 Results for $K=256$
In Figure 3, we observe the macro-F1 scores for models $K=256$ on the subfigures on the right. DPO-augmentation continues to enhance Baseline models which already benefits from more training data relative to lower $K$ values, with improvement in scores reaching up to 13.2% for Llama and 14% for T5. In contrast, KTO-augmentation appears ineffectual on HateXplain and even hinders performance on Latent Hate and Implicit Hate, showing notable reduction from Baseline scores, respectively. With this, we hypothesize that DPO’s contrastive nature reinforces signals that allow the LLM to learn from its own mistakes, thereby sharpening its reasoning ability for this type of discriminatory task. On the other hand, KTO’s reliance on implicit signals from listwise data is not sufficient to improve discernment of classes when the semantic distinction between labels are nuanced for Latent Hate and Implicit Hate.
@@
HateXplain Latent Hate Implicit Hate
Model F1 F1% Data% F1 F1% Data% F1 F1% Data%
SMARTER (Ours)
   Llama_DPO-256 0.64 89% 6% 0.69 100% 7% 0.60 90% 57%
   T5_DPO-256 0.62 86% 6% 0.65 94% 7% 0.59 88% 57%
Full Training Baselines
   Llama_Full 0.72 100% 100% 0.62 90% 100% 0.67 100% 100%
   T5_Full 0.72 100% 100% 0.61 88% 100% 0.64 96% 100%
   ModernBERT 0.70 97% 100% 0.61 88% 100% 0.64 96% 100%
Commercial: Zero-shot
   GPT-4o-mini 0.50 69% 0.54 78% 0.42 63%
   GPT-4.1 0.55 76% 0.60 87% 0.57 85%
   GPT-5-chat 0.56 78% 0.51 74% 0.58 87%
   Qwen-32B 0.54 75% 0.47 68% 0.49 73%
Commercial: 16-shot ICL
[…]
Ablation and Baseline Comparisons.
To isolate SMARTER’s contribution, Table 5 shows off-the-shelf models without HateCOT pretraining achieve 0.52 F1 on HateXplain. Adding HateCOT (Baseline $K=256$) improves to 0.58 (+0.06), while SMARTER’s DPO self-augmentation reaches 0.64 (+0.06 additional), contributing performance gain.
[…]
@@ 6 Stage 2: Cross-model Refinement
Having shown individual models self-improve efficiently, we explore whether cross-model training can further enhance performance by transferring explanation quality between LLMs. We first conduct human evaluation of both models’ explanations, then train each model on the other’s outputs.
@@
Table 2: Annotator preferences for explanations on 342 HateXplain samples (majority vote).
6.1 Human Evaluation
To simulate real-life deployment, we solicit 14 crowdworkers from diverse demographic backgrounds on the platform Amazon Mechanical Turks (AMT) to annotate their preference on the explanations offered by the LLMs in this study . Annotators are encouraged to evaluate the pairs of explanation for each post based on the criteria of Clarity, Reasoning and Alignment as illustrated with the template in Figure 5. To avoid position bias, we present the pairs of explanations in randomized orders. All annotators are compensated fairly.
@@ 6.2 Cross-model Refinement
Motivated by Llama’s superior human-rated explanations, we investigate whether models can benefit from each other’s outputs through cross-training. Below we describe our cross-model methodology, evaluate its impact on classification accuracy, and analyze how explanation styles evolve.
@@ 6.2.1 Cross-model Training Methodology
While Llama produces human-preferred explanations and is the natural deployment choice, we investigate whether models can adopt each other’s explanation styles without degrading classification performance. Using the prompting technique described in Section 6.1, we use the SFT+DPO $K=128$ variants of both Llama and T5 LLMs to generate explanations conditioned on the gold labels of the held-out 128-shot portion of the training data (unseen by either model). We then perform cross-model training: T5 is finetuned on Llama’s outputs and vice versa, applying both SFT and DPO self-augmentation as in Section 5.1.
@@ 6.2.3 Cross-model Style Analysis
Figure 12 shows individually trained models exhibit their own styles, while cross-model SFT aligns outputs with the alternative’s style. Except for T5 on HateXplain, cross-model DPO slightly amplifies the base model’s style, suggesting this mechanism induces mild stylistic convergence.
@@ 7 Analysis of Explanations
Using Natural Language Inference (NLI), we assess explanation consistency with labels and definitions, then analyze cross-model stylistic changes.
@@ 7.1 Analysis of Consistency
We prompt GPT-4o-mini with the template shown in Figure 9 to judge whether the offered explanations are logically consistent with the predicted label for each sample across 3 test sets via the 4 categories: Entail, Contradict, Neutral and Undefined (noise cases where either the predicted label or the explanation is missing). An author’s independent annotation of 150 random samples from this pool under the same instruction achieves nearly unanimous agreement with GPT-4o-mini (Cohen’s $\kappa=0.92$), an observation inline with recent works that validate LLM-as-a-judge approach in NLI applications Negru et al. (2025); Gallipoli and Cagliero (2025).
@@ 8 Discussion and Conclusion
Our empirical analysis yields several key recommendations to guide practitioners in deployment.
SMARTER delivers controllable, cost-efficient moderation.
Our staged pipeline achieves 86-100% of full-model performance using only 6-57% of training data (Table 1). Notably, we demonstrate this efficiency on challenging multi-class settings, whereas prior works tend to focus on simpler tasks Guo et al. (2023); Kim et al. (2024). Unlike commercial APIs—where 16-shot ICL exhibits high variance ($\pm$0.10) and can degrade performance (Table 6)—SMARTER produces deterministic, explainable outputs suitable for production deployment at a fraction of the cost. When labeled data is scarce and explainability is required, SMARTER provides cost-efficient, transparent production deployment versus commercial APIs.
Cross-model refinement enables targeted improvement.
T5’s superior cross-trained F1 scores (Figure 4) show that weaker models can adopt stronger models’ reasoning patterns. Practitioners facing category-specific weaknesses (Figure 11) can train multiple LLMs, then apply cross-model finetuning at high-quality checkpoints (e.g., $K=128$) to transfer strengths. Architectural diversity (encoder-decoder vs. decoder-only) appears beneficial, as Llama enhances reasoning diversity that T5 leverages Hao et al. (2024).
Human oversight remains essential.
Human evaluation (Table 2) shows category-specific preferences: Llama excels on Normal posts (3:1) while matching on Hate. Practitioners should periodically validate explanations, especially after cross-model training, to maintain consistency (Table 3).